Three resistors, 2.62, 3.86, and 4.28Ω, are connected in series across a 15.3-V battery. Find the power delivered to each resistor.?
1
Expert's answer
2020-09-09T10:21:37-0400

solution:-

given data

resistance(R1)="2.62 \Omega"

resistance(R2)=3.86"\Omega"

resistance(R3)=4.28"\Omega"

voltage(V)=15.3V


all the resister are in series so equivalent resistance can obtain as


"R=R_1+R_2+R_3"

by putting the value


"R=2.62+3.86+4.28=10.76\Omega"


all the resister are in series so current will be same in all resistance


current(I)="\frac{V}{R}"


by putting value of V and R


"I=\frac{15.3}{10.76}=1.42A"


power in any resister can be obtain as


"P=I^2R"


power in resistance R1


"P_1=1.42\times1.42\times2.62=5.28W"

power in resistance R2

"P_2=1.42\times1.42\times3.86=7.76W"


power in resistance R3

"P_3=1.42\times1.42\times4.28=8.60W"


so power in R1 ,R2 ,and R3 are 5.28W , 7.76W and 8.60W respectively.